# Stop treating ‘AGI’ as the north-star goal of AI research

### v3

@@ 4.1 Alternative View
Thoughtful attempts to address shortcomings in accounts of AGI indeed exist (e.g., Adams et al., 2012; Chollet, 2019; Morris et al., 2023; Summerfield, 2023). If prior accounts of AGI are counterproductive or flawed, why not pursue new accounts that address those flaws?
As an example of the alternative view, Morris et al. (2023) arguably mitigate the Illusion of Consensus trap by disentangling the disagreements about goals, predictions, and risks that plague prior accounts of AGI. Moreover, their proposal for a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021) could be viewed as mitigating the Presuming Value-Neutrality trap. Setting society-wide standards could be done in a way that explicitly centers specific risks that the standards address (Morris et al., 2023). In this way, the values being favored manifest in the risks that are centered by the standards. Such works showcase a potential response to traps. In arguing that AGI discourse aggravates multiple standing problems, we cannot rule out the possibility of efforts that mitigate these same problems while retaining AGI as a goal. Moreover, given that lack of agreement about how to define AGI is likely to persist, as Morris et al. (2023), Summerfield (2023), and many others do, it would be especially implausible for us to presume to have a complete enough view of the landscape of possible conceptions of AGI to draw definitive conclusions.
@@ 4.2 Rebuttal
Reason 1: Conflict with our recommendations. First, we believe that specificity of goals and pluralism of goals and approaches are at odds with this alternative view. Although the definition of AGI is highly contested, a frequent motivation for embracing AGI as a north-star goal is the desire for a single, large-scale, unifying vision for the field (Summerfield, 2023).
Reason 2: Distinguishing hype from reality. Another reason to favor our position is the research community’s responsibility to help distinguish hype from reality. We believe that the AI research community must find a path to providing trusted, evidence-based answers to increasingly complex questions about AI technologies, their goals, and their impacts. In the current moment, the hyped terminology of AGI tends to undermine this goal.
No matter how well or poorly defined, AGI has acquired a cultural significance that exacerbates the challenge of distinguishing hype from reality. “Intelligence” and “generality” hold the promise of being beneficial for countless needs and contexts (§2.32.5). No matter how cautious the research community attempts to be, the cultural associations of these terms risk stoking the flames of unscientific thinking about AI. This enables various interested parties to loosely project utopian or dystopian characteristics onto AGI in ways that support their calls for more power and resources.
Reason 3: Benefiting humans as the goal of technology. We worry that large-scale stories about unifying goals for AI are prone to blurring the line between hype and reality. If the AI community nevertheless wants an overarching goal to strive towards, the goal should be the support and benefit of human beings. The goals of technology are shaped by people. Evidence-based approaches to examining whether technology effectively meets the needs of people—be they “users”, “consumers”, “patients”, “scholars”, or a myriad of business and social monikers—are well-established. In a quest to achieve AGI, communities often lose sight of the needs of people as a goal, in favor of focusing on just the technology.
@@
5 Acknowledgments
Borhane Blili-Hamelin was funded in part through a Magic Grant from the Brown Institute for Media Innovation. Chris Graziul was supported by the National Institute of Minority Health and Health Disparities of the National Institutes of Health under award number R01MD015064. This material is based upon work supported in part by the NSF Program on Fairness in AI in Collaboration with Amazon under Award IIS-2147305. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, National Institutes of Health, the Brown Institute for Media Innovation, or Amazon.
@@ Appendix A Definitions of AGI and related concepts
Table 1: Sample definitions of AGI and related concepts.
Sample of proposed definitions of AGI and related concepts—human-level AI, strong AI, artificial superintelligence (ASI), etc.
Weizenbaum (1976). “Intelligence is a meaningless concept in and of itself. It requires a frame of reference, a specification of a domain of thought and action, in order to make it meaningful. […] [T]hese domains are themselves not measurable. […].” Argues that any argument that calls for the conclusion or denial that “machines may surpass us in general intelligence” is “ill-framed and therefore sterile” due to “our inability to compute an upper bound on machine intelligence”. We follow Blili-Hamelin et al. (2024) in considering this critical account relevant to debates about how to conceive AGI.
@@
Nilsson (2005). “achieving real human-level artificial intelligence would necessarily imply that most of the tasks that humans perform for pay could be automated. Rather than work toward this goal of automation by building special-purpose systems, I argue for the development of general-purpose, educable systems that can learn and be taught to perform any of the thousands of jobs that humans can perform.
Fast Company (2010)Could a computer make a cup of coffee?” Tasks the machine to go into an “average” American home, find ingredients, and make a cup of coffee. This requires embodied AI systems. Wozniak’s test has since been included in discussions of AGI. (Goertzel et al., 2012)
@@
Goertzel (2014). “roughly speaking, an AGI system is a synthetic intelligence that has a general scope and is good at generalization across various goals and contexts.”.
@@
Chollet et al. (2024); Chollet (2019). Defines AGI as“a system capable of efficiently acquiring new skills and solving novel problems for which it was neither explicitly designed nor trained”. In 2019, introduced an as yet (January 2025) unsolved benchmark for incentivizing progress towards AGI thus defined. Proposes that “it’s still feasible to create unsaturated, interesting benchmarks that are easy for humans, yet impossible for AI – without involving specialist knowledge. We will have AGI when creating such evals becomes outright impossible” (Chollet, 2024b).
@@
Hernández-Orallo et al. (2021). “independently of its overall capability, an agent can only be called fully general if it covers all tasks up to an equivalent level of difficulty, determined by the resources that are needed for them”. Introduces two individual-specific (be them humans, other animals or AI systems) measures that, together, decouple the concept of general intelligence: (i) ‘generality’, which refers to “[…] the distribution of the tasks the agent can solve,” and (ii) ‘capability’, which indicates “[…] how far, on average, an agent can reach in terms of task difficulty”.
@@
Marcus (2022). Defines AGI as “a shorthand for any intelligence … that is flexible and general, with resourcefulness and reliability comparable to (or beyond) human intelligence”. Calls for the need to operationalize the definition as a single system that can succeed in at least 3 of 5 proposed tasks, including Wozniak’s coffee cup benchmark (Fast Company, 2010).
@@
Morris et al. (2023). Proposes a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021). Describes a graded set of levels of achievement of target characteristics that can each be associated with tangible “metrics”, the introduction of “risks”, and changes in “Human-AI Interaction paradigm” (Morris et al., 2023). The framework targets 2 characteristics: levels of performance (which they define as “the depth of an AI system’s capabilities, i.e., how it compares to human-level performance for a given task”), and levels of generality (defined as “breadth of an AI system’s capabilities, i.e., the range of tasks for which an AI system reaches a target performance threshold.”).
@@
Summerfield (2023) We consider this account to endorse a strong form of pluralism about intelligence: natural intelligence takes many different shapes across cultures and species, serving many goals and functions that are irreducibly shaped by “the internal model by which an animal understands the world”, which itself “depends on its local environment, its embodied form, its desires and goals, and its interactions with conspecifics.” The same should be expected for “strong AI” or “AGI”. However, building on Dreyfus & Dreyfus (1986), Summerfield argues we have practical reasons to constrain the forms AI takes. The goal of AI is “to help humans in their endeavours.” To that end,“if we want to build AI systems that exhibit human-like intelligence, with whom we can interact in pursuit of human-centred goals, these agents will need to think in ways that make sense to us,”.777“[W]e are building AI to make the world a better place. But if we want AI to be useful to people, it will need to share our umwelt. If we build an AI that sees the world in a radically different way to us, its behaviour and mental states will be unintelligible. Such an agent will be at best unreliable and at worst unsafe.” (Summerfield, 2023)
@@
Attard-Frost (2023) Defines human and artificial intelligence as “value-dependent cognitive performance”, and “centres interdependencies between agents, their environments, and their measurers in collectively constructing and measuring context-specific performances of intelligent action”. Although this account is not presented as a conception of AGI, Blili-Hamelin et al. (2024) argue that the account is relevant to the topic.
@@
Suleyman & Bhaskar (2023)Artificial intelligence (AI) is the science of teaching machines to learn humanlike capabilities. Artificial general intelligence (AGI) is the point at which an AI can perform all human cognitive skills better than the smartest humans.

### v4

@@ 4.1 Counterargument
Thoughtful attempts to address shortcomings in accounts of AGI indeed exist (e.g., Adams et al., 2012; Chollet, 2019; Summerfield, 2023; Morris et al., 2024). If prior accounts of AGI are counterproductive or flawed, why not pursue new accounts that address those flaws?
As an example of an alternative view, Morris et al. (2024) arguably mitigate the Illusion of Consensus trap by disentangling the disagreements about goals, predictions, and risks that plague other accounts of AGI. Moreover, their proposal for a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021) could be viewed as mitigating the Presuming Value-Neutrality trap. Setting society-wide standards could be done in a way that explicitly centers specific risks that the standards address (Morris et al., 2024). In this way, the values being favored manifest in the risks that are centered by the standards. Such works showcase a potential response to traps. In arguing that AGI discourse aggravates multiple standing problems, we cannot rule out the possibility of efforts that mitigate these same problems while retaining AGI as a goal. Moreover, given that lack of agreement about how to define AGI is likely to persist, as Summerfield (2023), Morris et al. (2024), and many others believe, it would be especially implausible for us to presume to have a complete enough view of the landscape of possible conceptions of AGI to draw definitive conclusions.
@@ 4.2 Rebuttal
Reason 1: Conflict with our recommendations. Although the definition of AGI is highly contested, a frequent motivation for embracing AGI as a north-star goal is the desire for a single, large-scale, unifying vision for the field (Summerfield, 2023). This is somewhat in tension with our recommendation of goal pluralism, which we argue is valuable for AI research. “AGI” also brings with it a notion of “general” that can be discordant with our recommendation of specificity. [^9] Note that these are not all things considered (or definitive) reasons to abandon the alternative view. Rather, readers who are convinced by our arguments in favor of specificity and pluralism have, to some extent, reasons to be wary of the project of improving AGI. These are so-called pro-tanto (i.e., “to that extent”) reasons, which can be overridden by other considerations (Alvarez, 2023). As such, there are reasons to be wary of AGI-focused alternatives.
Reason 2: Distinguishing hype from reality. Another reason to favor our position is the AI research community’s responsibility to help distinguish hype from reality. We believe that our community must provide trusted, evidence-based answers to increasingly complex questions about AI technologies, their goals, and their impacts. In the current moment, the hyped terminology of AGI undermines this responsibility.
No matter how well or poorly defined, AGI has acquired a cultural significance that exacerbates the challenge of distinguishing hype from reality. “Intelligence” and “generality” hold the promise of being beneficial for countless needs and contexts (§2.3, §2.5). No matter how cautious the research community attempts to be, the cultural associations of these terms risk stoking the flames of unscientific thinking about AI. This enables various parties to loosely project utopian or dystopian characteristics onto AGI in ways that support their calls for more power and resources.
Reason 3: Benefiting humans as the goal of technology. If the AI community nevertheless wants an overarching goal to strive towards, the goal should be the support and benefit of human beings. The goals of technology are shaped by people. Evidence-based approaches to examining whether technology effectively meets the needs of people—be they “users”, “consumers”, “patients”, “scholars”, or a myriad of business and social monikers—are well-established. In a quest to achieve AGI, communities often lose sight of the needs of people as a goal, in favor of focusing on just the technology.
@@
Acknowledgments
Borhane Blili-Hamelin was funded in part through a Magic Grant from the Brown Institute for Media Innovation. Christopher Graziul was supported by the National Institute of Minority Health and Health Disparities of the National Institutes of Health under award number R01MD015064. This material is based upon work supported in part by the NSF Program on Fairness in AI in Collaboration with Amazon under Award IIS-2147305. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, National Institutes of Health, the Brown Institute for Media Innovation, or Amazon.
@@ Appendix A Definitions of AGI and Related Concepts
Appendix B AGI as a North-Star Goal
This paper argues against AGI serving as a north-star goal of AI research. When we talk about AGI being treated as north-star goal, we are not claiming that the majority of AI researchers are explicitly working in pursuit of this goal. In fact, many AI researchers may, like us, doubt or reject this goal (Salvaggio, 2025; Francesca Rossi et al., 2025). Rather, we claim that influential researchers and executives hold this view, enough so for it to deserve scrutiny. These dominant voices are further amplified by the publicity that discussions of AGI generate, including by members of the press (Klein, 2025), and by government commissions (Bommasani et al., 2025). As such, AGI has come to permeate both community incentives and cultural norms. In this context, interrogating its role and influence in the AI research community matters. Below we provide a small sample of quotes and resources illustrating this effect.
OpenAI
The mission statement of OpenAI is “…to ensure that [AGI] benefits all of humanity.” Its website (OpenAI, 2025a) states that “we are building safe and beneficial AGI, but will also consider our mission fulfilled if our work aids others to achieve this outcome.” This goal directly influences both the company’s direct work (OpenAI, 2025b) and the work that it funds (OpenAI, 2025c).
Google DeepMind
The vision statement (DeepMind, 2025) of Google DeepMind states that “[AGI] has the potential to drive one of the greatest transformations in history.” In a recent briefing (Browne, 2025), Demis Hassabis stated that, though current systems still have limitations, over the next 5–10 years “a lot of those capabilities will start coming to the fore and we’ll start moving towards what we call [AGI].” A position paper (Morris et al., 2024) by Google DeepMind authors last year defines concrete goals in pursuit of AGI.
Anthropic
In the essay “Machines of Loving Grace,” Anthropic CEO Dario Amodei argues how the world could be shaped positively by “Powerful AI” aligned with “AGI” goals  (Amodei, 2024). Amodei recently told CNBC that AI that is “better than almost all humans at almost all tasks” can emerge shortly (Browne, 2025). Anthropic’s official recommendations (Anthropic, 2025) to OSTP for the U.S. AI Action Plan state that “we expect powerful AI systems will emerge [with] intellectual capabilities matching or exceeding that of Nobel Prize winners.”
Other Influential Executives & Researchers
In “Sparks of AGI” (Bubeck et al., 2023), researchers at Microsoft argue that GPT-4 “could reasonably be viewed as an early (yet still incomplete) version of [AGI].” In announcing xAI, Elon Musk stated (Kerner, 2023) that “the overarching goal of xAI is to build a good AGI.” Speaking to TIME (Perrigo, 2024), Yann LeCunn explained that he refers to “what people call ‘AGI”’ as “human-level intelligence,” and noted that “the mission of FAIR [Meta’s Fundamental AI Research team] is human-level intelligence.” Geoff Hinton stated to Forbes (Maymin, 2023) that he “is certain we will have AGI soon, and biological humans will be relegated to be the second-smartest species on the planet.”
Table 1: Sample of proposed definitions of AGI and related concepts.
Weizenbaum (1976). “Intelligence is a meaningless concept in and of itself. It requires a frame of reference, a specification of a domain of thought and action, in order to make it meaningful. […] [T]hese domains are themselves not measurable.” Argues that any argument that calls for the conclusion or denial that “machines may surpass us in general intelligence” is “ill-framed and therefore sterile” due to “our inability to compute an upper bound on machine intelligence.” We follow Blili-Hamelin et al. (2024) in considering this critical account relevant to debates about how to conceive AGI.
@@ Appendix B AGI as a North-Star Goal
Nilsson (2005). “achieving real human-level artificial intelligence would necessarily imply that most of the tasks that humans perform for pay could be automated. Rather than work toward this goal of automation by building special-purpose systems, I argue for the development of general-purpose, educable systems that can learn and be taught to perform any of the thousands of jobs that humans can perform.”
Fast Company (2010)Wozniak: Could a Computer Make a Cup of Coffee?” tasks the machine to go into an “average” American home, find ingredients, and make a cup of coffee. This requires embodied AI systems. Wozniak’s test has since been included in discussions of AGI (Goertzel et al., 2012).
@@
Goertzel (2014). “roughly speaking, an AGI system is a synthetic intelligence that has a general scope and is good at generalization across various goals and contexts.”
@@
Chollet (2019); Chollet et al. (2024). Defines AGI as “a system capable of efficiently acquiring new skills and solving novel problems for which it was neither explicitly designed nor trained.” In 2019, introduced an as yet (January 2025) unsolved benchmark for incentivizing progress towards AGI thus defined. Proposes that “it’s still feasible to create unsaturated, interesting benchmarks that are easy for humans, yet impossible for AI – without involving specialist knowledge. We will have AGI when creating such evals becomes outright impossible” (Chollet, 2024b).
@@
Hernández-Orallo et al. (2021). “independently of its overall capability, an agent can only be called fully general if it covers all tasks up to an equivalent level of difficulty, determined by the resources that are needed for them.” Introduces two individual-specific (be them humans, other animals or AI systems) measures that, together, decouple the concept of general intelligence: (i) ‘generality’, which refers to “the distribution of the tasks the agent can solve,” and (ii) ‘capability’, which indicates “how far, on average, an agent can reach in terms of task difficulty.”
@@
Marcus (2022). Defines AGI as “a shorthand for any intelligence… that is flexible and general, with resourcefulness and reliability comparable to (or beyond) human intelligence.” Calls for the need to operationalize the definition as a single system that can succeed in at least 3 of 5 proposed tasks, including Wozniak’s coffee cup benchmark (Fast Company, 2010).
@@
Morris et al. (2024). Proposes a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021). Describes a graded set of levels of achievement of target characteristics that can each be associated with tangible “metrics”, the introduction of “risks”, and changes in “Human-AI Interaction paradigm” (Morris et al., 2024). The framework targets 2 characteristics: levels of performance (which they define as “the depth of an AI system’s capabilities, i.e., how it compares to human-level performance for a given task”), and levels of generality (defined as “breadth of an AI system’s capabilities, i.e., the range of tasks for which an AI system reaches a target performance threshold”).
@@
Summerfield (2023) We consider this account to endorse a strong form of pluralism about intelligence: natural intelligence takes many different shapes across cultures and species, serving many goals and functions that are irreducibly shaped by “the internal model by which an animal understands the world”, which itself “depends on its local environment, its embodied form, its desires and goals, and its interactions with conspecifics.” The same should be expected for “strong AI” or “AGI”. However, building on Dreyfus & Dreyfus (1986), Summerfield argues we have practical reasons to constrain the forms AI takes. The goal of AI is “to help humans in their endeavours.” To that end,“if we want to build AI systems that exhibit human-like intelligence, with whom we can interact in pursuit of human-centred goals, these agents will need to think in ways that make sense to us.”101010“[W]e are building AI to make the world a better place. But if we want AI to be useful to people, it will need to share our umwelt. If we build an AI that sees the world in a radically different way to us, its behaviour and mental states will be unintelligible. Such an agent will be at best unreliable and at worst unsafe.” (Summerfield, 2023)
@@
Attard-Frost (2023) Defines human and artificial intelligence as “value-dependent cognitive performance”, and “centres interdependencies between agents, their environments, and their measurers in collectively constructing and measuring context-specific performances of intelligent action.” Although this account is not presented as a conception of AGI, Blili-Hamelin et al. (2024) argue that the account is relevant to the topic.
@@
Suleyman & Bhaskar (2023) “Artificial intelligence (AI) is the science of teaching machines to learn humanlike capabilities. Artificial general intelligence (AGI) is the point at which an AI can perform all human cognitive skills better than the smartest humans.”
@@
Author Contributions
We follow the CRediT recommendations and taxonomy provided by Allen et al. (2019) to determine and outline author contributions. [^11] The initial ICML submission was created by Borhane Blili-Hamelin, Leif Hancox-Li, and Christopher Graziul, leveraging extensive work and writing from the larger project. All contributors then worked together on refining this submission.
  • Borhane Blili-Hamelin: Conceptualization (Formulation & Evolution), Investigation, Methodology (Development), Project administration, Supervision (Oversight & Leadership), Writing (Initial draft, Submitted draft, Review & Editing).
  • Christopher Graziul: Conceptualization (Formulation & Evolution), Investigation, Methodology (Development), Project administration, Supervision (Oversight & Leadership), Writing (Initial draft, Submitted draft, Review & Editing).
  • Leif Hancox-Li: Conceptualization (Formulation & Ideas), Investigation, Methodology (Development), Supervision (Mentorship), Writing (Initial draft, Submitted draft, Review & Editing).
  • Hananel Hazan: Conceptualization (Ideas &  Evolution), Investigation, Methodology (Development), Project administration, Writing (Initial draft, Review & Editing, ).
  • El-Mahdi El-Mhamdi: Conceptualization (Evolution & Ideas), Writing (Initial draft, Review & Editing).
  • Avijit Ghosh: Conceptualization (Ideas), Writing (Submitted draft [Traps section], Review & Editing).
  • Katherine Heller: Conceptualization (Formulation & Evolution), Supervision (Mentorship), Writing (Initial draft, Submitted draft, Review & Editing).
  • Jacob Metcalf: Conceptualization (Formulation & Evolution), Writing (Submitted draft, Review & Editing).
  • Fabricio Murai: Conceptualization (Ideas), Writing (Submitted draft [Table of AGI definitions], Review & Editing, ).
  • Eryk Salvaggio: Conceptualization (Evolution & Ideas), Writing (Submitted draft [Introduction, Traps section], Review & Editing).
  • Andrew Smart: Conceptualization (Formulation & Evolution), Writing (Initial draft, Review & Editing).
  • Todd Snider: Conceptualization (Ideas), Methodology (Implementation), Writing (Initial draft, Review & Editing, ).
  • Mariame Tighanimine: Conceptualization (Evolution & Ideas), Writing (Initial draft, Review & Editing).
  • Talia Ringer: Conceptualization (Formulation, & Ideas), Project administration, Supervision (Oversight & Leadership), Writing (Initial draft, Review & Editing).
  • Margaret Mitchell: Conceptualization (Ideas &  Evolution), Investigation, Methodology (Development), Supervision (Oversight & Mentorship), Writing (Initial draft, Submitted draft [all sections], Review & Editing, Rebuttal).
  • Shiri Dori-Hacohen: Conceptualization (Ideas & Evolution), Investigation, Methodology, Project administration, Supervision (Oversight & Leadership), Writing (Initial draft, Submitted draft [all sections], Review & Editing).